the width of a box is 4 inches more then the height. the length is the difference of 9 inches and the height. the volume of the box is 180 cubic inches. what are all the poswible dimensions for the box?
Answer by Alan3354(69443)

the width of a box is 4 inches more then the height. the length is the difference of 9 inches and the height. the volume of the box is 180 cubic inches. what are all the poswible dimensions for the box?
-------------
W = H + 4
L = 9 - H
---
Vol = L*W*H = H*(H+4)*(9-H) = 180
---
H^3 - 5H^2 - 36H + 180 = 0
(H-5)*(H-6)*(H+6) = 0
H = 5 or 6
--> 5 by 9 by 4
or 6 by 10 by 3